Circular No. 6614

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) D. Mehringer, Caltech Submillimeter Observatory (CSO); P. Colom,

Observatoire de Paris (OP); D. Benford, CSO; D. Bockelée-Morvan, OP; D. Despois, Observatoire de Bordeaux; G. Paubert, Institut de Radioastronomie Millimétrique (IRAM); B. Germain, N. Biver, J. Crovisier, D. Gautier, E. Gérard, and H. Rauer, OP; D. C. Lis and T. G. Phillips, CSO; R. Moreno, IRAM; and J. K. Davies and W. R. F. Dent, Joint Astronomy Centre, Hilo, report the identification of formamide (NH_{2}CHO) in comet C/1995 O1: “On Apr. 5 UT, we detected the lines of NH_{2}CHO at 254.877 and 227.606 GHz with the CSO and IRAM 30-m telescopes, respectively. The integrated line areas in the main-beam brightness temperature scale are 0.070 ± 0.009 (254.9 GHz) and 0.28 ± 0.02 (227.6 GHz) K km/s. The line at 218.459 GHz was also marginally detected (0.26 ± 0.06 K km/s) with the IRAM 30-m telescope. A preliminary estimate of the NH_{2}CHO production rate is 3 to 5 × 10^{27} molecules/s. The 252-GHz lines of methanol observed with the CSO on Apr. 3 agree with a rotational temperature of 119 ± 9 K.”

A. Owens, T. Oosterbroek, A. Orr, and A. N. Parmar, Space Science Department, European Space Agency; L. A. Antonelli and F. Fiore, BeppoSAX Science Data Center, Rome; M. C. Maccarone, Istituto di Fisica Cosmica e Applicazioni Informatica, CNR, Palermo; and L. Piro, Insituto Astrofisica Spaziale, CNR, Frascati, on behalf of the BeppoSAX team, report the detection of soft x-rays from the direction of comet C/1995 O1 by the LECS instrument onboard BeppoSAX: “The observations took place between 1996 Sept. 10.16 and 11.18 UT. After correcting for the comet’s motion, a 10σ enhancement was found within 3′ of the expected location of the nucleus at much greater than 99.99-percent confidence. A weak, 2.7σ enhancement was also found in motion-corrected MECS data at the same location. The extracted LECS spectrum is well fit by a thermal bremsstrahlung model of kT = 0.36 keV — consistent with that observed in other comets (e.g., Lisse et al. 1996, Science 274, 205). The total 0.10–2.0keV luminosity is 6 × 10^{16} erg/s, comparable with the observed luminosities of comets C/1996 B2 and C/1990 N1. There is no evidence for fluorescent carbon or oxygen emission with 95-percent-confidence upper flux limits of 1 × 10^{−4} photons cm^{−2} s^{−1} and 3 × 10^{−4} photons cm^{−2} s^{−1} for narrow line emission at 282 and 523 eV, respectively. This imples that if lines are present, their total luminosity must be < 10 percent of the continuum luminosity.”

1997 April 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green